Problem 26.91

The index of refraction for red light in a certain liquid is 1.322; the index of refraction for violet light in the same liquid is 1.333.

Part A

Find the dispersion ?v??r for red and violet light when both are incident on the flat surface of the liquid at an angle of 40.00? to the normal.

Express your answer using two significant figures.

?v??r =   ?  

SubmitMy AnswersGive Up
Answer #1

use snells law

sin(theta_i)/sin(theta_r) = n2/n1


for red light

sin(40)/sin(theta_R) = 1.322/1

==> theta_R = sin^-1(sin(40)/1.322)

= 29.09 degrees


for volilet light

sin(40)/sin(theta_V) = 1.333/1

==> theta_V = sin^-1(sin(40)/1.333)

= 28.83 degrees

theta_V - Theta_R = 28.83 - 29.09

= -0.26 degrees
Answer #2

inverse sin (1*sin(40))/1.333 - inverse sin (1*sin(40))/1.322

=arc sin (0.4822) - arc sin(0.4862) =-0.2619 degree

The dispersion for two significant figures is 0.26 degree